A monkey of mass 20kg is holding a vertical rope.The rope will not break when a mass of 25 kg is suspended on it,but will break when the mass exceed 25 kg.w hat is the maximum acceleration with which the monkey can climb up along the road.

Grade:

1 Answers

Susmita
5 years ago
F=ma
Force here is 25 kg.w=25×g=25×9.81 Newton
Mass m=20 kg
Acceleration  a=F/m = (25×9.81)/20 =12.26 m/s^2